OpenStudy (anonymous):

M_1N_1=M_2N_2O solve for N_2

OpenStudy (anonymous):

that question looks confusing

OpenStudy (anonymous):

\[m _{1}n _{1}=m _{2}n _{2}\]

OpenStudy (shane_b):

Just divide both sides by N_1

OpenStudy (shane_b):

\[m _{1}=\frac{m _{2}n _{2}}{n _{1}}\]

OpenStudy (anonymous):

ohh ok thx

OpenStudy (anonymous):

how would i get solve for n2

OpenStudy (shane_b):

Same procedure...just divide both sides of the the original equation by m_2 instead (so that m_2 moves to the other side and n_2 is isolated).

OpenStudy (anonymous):

would i divide the O also

OpenStudy (anonymous):

i forgot to add the O

OpenStudy (anonymous):

the O is after N2

OpenStudy (shane_b):

Same deal...you would just divide both sides by m_2O. If the original equation was actually this:\[m_1n_1=m_2n_2O\]Solving it for m1 would be:\[m_1=\frac{m_2n_2O}{n_1}\]Solving it for n2 would be:\[n_2=\frac{m_1n_1}{m_2O}\]